# Non-analyticity of the $S$-matrix

### v1

@@ 3 LSZ reduction formula and lack of analyticity
In this section we will derive the LSZ reduction formula [14], following Weinberg’s “polology” arguments [15]. This is well-known material in the Lorentz invariant case, and we will highlight the differences in the absence of Lorentz invariance.
@@
$\mathcal{T}=i\int\mathrm{d}^{4}z\,e^{iq\cdot z}\langle\boldsymbol{p}_{2}|\theta(z^{0})\left[\frac{\partial_{z^{0}/2}^{2}+E_{-}(-i\partial_{z_{i}/2})^{2}}{Z_{-}^{\pi}(-i\partial_{z_{i}/2})}\pi\left(\frac{z}{2}\right),\frac{\partial_{z^{0}/2}^{2}+E_{-}(-i\partial_{z_{i}/2})^{2}}{\bar{Z}_{-}^{\pi}(-i\partial_{z_{i}/2})}\pi\left(-\frac{z}{2}\right)\right]|\boldsymbol{p}_{1}\rangle\,.$ (44)
In the Lorentz invariant case the $Z$’s are just constants and the differential operators are the Klein-Gordon ones, $\Box+m^{2}$. Since these operators are local, the integrand is nonzero only for $z^{\mu}$ in the forward lightcone. Therefore $\mathcal{T}(q^{\mu})$ is analytic if $\mathrm{Im}\;q^{\mu}$ lies in the forward lightcone. (In this region all the derivatives of $\mathcal{T}$ are convergent integrals, assuming polynomial boundedness of the matrix elements, and it thus defines an analytic function.) This primitive domain of analyticity is the stepping stone for the use of dispersive arguments in the $S$-matrix [17, 16]. Here, however, the differential operators acting on the fields are non-local. In particular, from the explicit expressions of $Z_{-}^{\pi}$ and $E_{-}^{2}$ one sees the presence of branch points for complex spatial momenta. These singularities are present for any value of the time components of the momenta. This means there is no cone of analyticity for $\mathrm{Im}\;q^{\mu}$, contrary to the Lorentz invariant case. The only guaranteed region of analyticity, which is universal and independent of the model at hand, is for $\mathrm{Im}\;q^{0}>0$: this is just a consequence of the fact the integrand above is retarded, i.e. different from zero only for $z^{0}\geq 0$.
@@ 4.1.2 High energies
Another interesting limit to explore is the “high-energy limit”, when the energies of the scattered states are much larger than all the other scales of the problem. For simplicity we specialize to a specific kinematical configuration and set
$q_{1}^{0}=q_{2}^{0}=p_{1}^{0}=p_{2}^{0}\equiv\omega\,.$ (60)
The high-energy limit is therefore realized by
@@
$\mathcal{M}_{c}=$ $-\frac{\omega^{2}(\hat{p_{1}}\cdot\hat{q_{1}}+\hat{p_{1}}\cdot\hat{q_{2}}+\hat{q_{1}}\cdot\hat{q_{2}}-3)}{v^{2}}-\frac{3\mu^{2}\omega}{v^{3}}+$ (63)
@@
$\mathcal{M}_{s}=$ $\frac{5\omega^{2}(\hat{p_{1}}\cdot\hat{q_{1}}-1)}{2v^{2}}+\frac{\omega\left(6\mu^{4}+M^{2}v^{2}\hat{p_{1}}\cdot\hat{q_{1}}-M^{2}v^{2}\right)}{2\mu^{2}v^{3}}+$ (64)
@@
$\mathcal{M}_{t}=$ $\frac{\omega^{2}(1-\hat{q_{1}}\cdot\hat{q_{2}})}{2v^{2}}+\frac{M^{2}\omega(1-\hat{q_{1}}\cdot\hat{q_{2}})}{2\mu^{2}v}+$ (65)
@@
$\mathcal{M}_{u}=$ $\frac{\omega^{2}(1-\hat{p_{1}}\cdot\hat{q_{2}})}{2v^{2}}+\frac{M^{2}\omega(1-\hat{p_{1}}\cdot\hat{q_{2}})}{2\mu^{2}v}+$ (66)
@@
$\mathcal{M}=-2\frac{M^{2}}{v^{2}}+\mathcal{O}\left(\frac{\Omega}{\omega}\right)=-4\lambda+\mathcal{O}\left(\frac{\Omega}{\omega}\right)~{}~{}\text{as }\frac{\omega}{\Omega}\rightarrow\infty\,.$ (67)
This constant value is equal to the $s\to\infty$ limit of the Lorentz invariant amplitude in Eq. (59): as expected, the high-energy limit of the amplitude reproduces its Lorentz invariant counterpart.
@@ Acknowledgements
It is a pleasure to thank B. Salehian for useful discussions.
@@ Appendix D Estimate of the decay width
$Z^{\pi}_{-}(\omega)=Z^{h}_{+}(\omega)\sim 1+\mathcal{O}\left(\frac{\omega^{2}}{\omega_{c}^{2}}\right)\>,\>Z^{\pi}_{+}(\omega)=Z^{h}_{-}(\omega)\sim-i\frac{\omega}{\omega_{p}}+\mathcal{O}\left(\frac{\omega^{3}}{\omega_{c}^{3}}\right)\,,$ (131)
@@
$\frac{\Gamma}{\omega}(\omega\gg\omega_{p})\sim\frac{\lambda^{2}v^{2}}{\omega^{2}}\sim\lambda\frac{M^{2}}{\omega^{2}}\ll\lambda\,,$ (139)

### v2

@@ 3 LSZ reduction formula and lack of analyticity
In this section we will derive the LSZ reduction formula [14], following Weinberg’s “polology” arguments [15]. [^5] We acknowledge Riccardo Rattazzi for clarifying certain aspects of Weinberg’s “polology” arguments and their link to the LSZ reduction formula, during the “Advanced Quantum Field Theory” course given at EPFL in 2020. This is well-known material in the Lorentz invariant case, and we will highlight the differences in the absence of Lorentz invariance.
@@
$\mathcal{M}=i\int\mathrm{d}^{4}z\,e^{iq\cdot z}\langle\boldsymbol{p}_{2}|\theta(z^{0})\left[\frac{\partial_{z^{0}/2}^{2}+E_{-}(-i\partial_{z_{i}/2})^{2}}{Z_{-}^{\pi}(-i\partial_{z_{i}/2})}\pi\left(\frac{z}{2}\right),\frac{\partial_{z^{0}/2}^{2}+E_{-}(-i\partial_{z_{i}/2})^{2}}{\bar{Z}_{-}^{\pi}(-i\partial_{z_{i}/2})}\pi\left(-\frac{z}{2}\right)\right]|\boldsymbol{p}_{1}\rangle\,.$ (44)
In the Lorentz invariant case the $Z$’s are just constants and the differential operators are the Klein-Gordon ones, $\Box+m^{2}$. Since these operators are local, the integrand is nonzero only for $z^{\mu}$ in the forward lightcone. Therefore $\mathcal{M}(q^{\mu})$ is analytic if $\mathrm{Im}\;q^{\mu}$ lies in the forward lightcone. (In this region all the derivatives of $\mathcal{M}$ are convergent integrals, assuming polynomial boundedness of the matrix elements, and it thus defines an analytic function.) This primitive domain of analyticity is the stepping stone for the use of dispersive arguments in the $S$-matrix [17, 16]. Here, however, the differential operators acting on the fields are non-local. In particular, from the explicit expressions of $Z_{-}^{\pi}$ and $E_{-}^{2}$ one sees the presence of branch points for complex spatial momenta. These singularities are present for any value of the time components of the momenta. This means there is no cone of analyticity for $\mathrm{Im}\;q^{\mu}$, contrary to the Lorentz invariant case. The only guaranteed region of analyticity, which is universal and independent of the model at hand, is for $\mathrm{Im}\;q^{0}>0$: this is just a consequence of the fact the integrand above is retarded, i.e. different from zero only for $z^{0}\geq 0$.
@@ 4.1.2 High energies
Another interesting limit to explore is the “high-energy limit”, intended as the limit achieved when the energies of the scattered states are much larger than all the other scales involved in the problem. In order to take this limit, for convenience, we specialize to the kinematical configuration in which the energies of the scattering states are all the same
$q_{1}^{0}=q_{2}^{0}=p_{1}^{0}=p_{2}^{0}\equiv\omega\,,$ (60)
and in which the angle between $\vec{q}_{1}$ and $\vec{p}_{1}$ is not zero. The high-energy limit is therefore realized by
@@
$\mathcal{M}_{c}=$ $-\frac{\omega^{2}(\hat{p_{1}}\cdot\hat{q_{1}}+\hat{p_{1}}\cdot\hat{q_{2}}+\hat{q_{1}}\cdot\hat{q_{2}}-3)}{v^{2}}-\frac{3\mu^{2}\omega}{v^{3}}$ (63)
@@
$\mathcal{M}_{s}=$ $\frac{5\omega^{2}(\hat{p_{1}}\cdot\hat{q_{1}}-1)}{2v^{2}}+\frac{\omega\left(6\mu^{4}+M^{2}v^{2}\hat{p_{1}}\cdot\hat{q_{1}}-M^{2}v^{2}\right)}{2\mu^{2}v^{3}}$ (64)
@@
$\mathcal{M}_{t}=$ $\frac{\omega^{2}(1-\hat{q_{1}}\cdot\hat{q_{2}})}{2v^{2}}+\frac{M^{2}\omega(1-\hat{q_{1}}\cdot\hat{q_{2}})}{2\mu^{2}v}$ (65)
@@
$\mathcal{M}_{u}=$ $\frac{\omega^{2}(1-\hat{p_{1}}\cdot\hat{q_{2}})}{2v^{2}}+\frac{M^{2}\omega(1-\hat{p_{1}}\cdot\hat{q_{2}})}{2\mu^{2}v}$ (66)
@@
$\mathcal{M}=-2\frac{M^{2}}{v^{2}}+\mathcal{O}\left(\frac{\Omega}{\omega}\right)=-4\lambda+\mathcal{O}\left(\frac{\Omega}{\omega}\right)~{}~{}\text{as }\>\frac{\omega}{\Omega}\rightarrow\infty\,.$ (67)
Interestingly, the amplitude asymptotes to the Lorentz-invariant amplitude for the scattering of $U(1)$ eigenstates (quanta of $\Phi$) in the unbroken phase and not to the $s\to+\infty$ limit of the Lorentz-invariant amplitude that describes the finite-angle scattering of Goldstones in the broken phase [^8] We thank the authors of [18] for pointing out a mistake in the first version of our paper.. Indeed, making use of (1), it is possible to check that the $\Phi\Phi\to\Phi\Phi$ scattering amplitude is given by
$\mathcal{M}_{\Phi\Phi\to\Phi\Phi}=-4\lambda\;,$ (68)
in agreement with (67). (On the other hand, the Lorentz-invariant amplitude of the Goldstones, Eq. (59), asymptotes to $-6\lambda$ in the $s\to+\infty$ limit at finite angle.) As discussed at the end of section 2.1, for $\omega\gg\omega_{c}$ the quanta of $\Phi$ are indeed the correct eigenstates. We refer the reader to reference [18] for a more comprehensive treatment and explanation.
@@ 5.2 $\omega\gg M$
It is important to stress that the non-analyticities studied in this paper are not a consequence of the absence of absolutely stable asymptotic states. Indeed, one can send the coupling $\lambda\to 0$ while keeping $\mu^{2}/v$ fixed. In this limit, $\Gamma\to 0$ while the discontinuity in $\mathcal{M}/\lambda$ remains constant.
@@ Acknowledgements
It is a pleasure to thank B. Salehian for useful discussions. MD and LS are partially supported by the SNSF grant 200021_213120.
@@ Appendix D Estimate of the decay width
$Z^{\pi}_{-}(\omega)=Z^{h}_{+}(\omega)\sim 1+\mathcal{O}\left(\frac{\omega^{2}}{\omega_{c}^{2}}\right)\>,\>Z^{\pi}_{+}(\omega)=Z^{h}_{-}(\omega)\sim-i\frac{\omega}{\omega_{c}}+\mathcal{O}\left(\frac{\omega^{3}}{\omega_{c}^{3}}\right)\,,$ (132)
@@
$\frac{\Gamma}{\omega}(\omega\gg\omega_{c})\sim\frac{\lambda^{2}v^{2}}{\omega^{2}}\sim\lambda\frac{M^{2}}{\omega^{2}}\ll\lambda\,,$ (140)